### v2

@@ 2.1 Post-trained Language Models
MODEL INPUT
[…]
Figure 3: Example of prompt Grattafiori et al. (2024). We omitted the end-of-turn token. The prompt always begins with a system and a user message. As shown, the system prompt includes the available tools as a JSON schema. Additionally, we also specify the function calling format for using the tools. Each call to an LM requires appending an assistant header for completion (not shown here). After execution of the tools, we append the tool response, and the assistant continues with the next completion.
From a probabilistic perspective, given an input $\mathbf{x}=(x_{1},\dots,x_{T})$ (also called a prompt) consisting of a sequence of tokens $x_{t}\in\mathcal{V}$, a post-trained LM defines a distribution over output sequences $\mathbf{y}=(y_{1},\dots,y_{N})$ conditioned on $\mathbf{x}$. The LM assistant generates $\mathbf{y}$ in an autoregressive manner:
@@
Post-training defines a broad objective where the LM is trained to be helpful, harmless, and honest Grattafiori et al. (2024). Typically, we use a mixture of supervised finetuning and preference learning Grattafiori et al. (2024). Notably, the prompt $\mathbf{x}$ consists of tasks such as function calling and instruction following.
As shown in Fig. 3, the input $\mathbf{x}$ corresponds to a chat where the sources of each message (typically, system, user, assistant, and tool) are specified by special headers. At inference, a user provides a query $\mathbf{q}$ and we parse it into a chat $\mathbf{x}$ with a system message and the query, optionally with a pre-appended user input template. In principle, at each step, the chat is updated with new messages from the assistant or tool. The system message contains, for instance, the general instructions, safety guidelines, constraints, and tools available. Additionally, post-trained LMs use function calling to interact with tools in principled ways. In formula, we denote a scaffolded LM as $\pi_{\phi}(\mathbf{q})$, where $\phi$ refers to non-parametric variables, including the system prompt and the user input template for parsing $\mathbf{q}$ into $\mathbf{x}$.
@@ 2.2 Tools
We follow the definition of Wang et al. Wang et al. (2024e) and define an LM-used tool as a function interface to a computer program that runs externally to the LM, where the LM generates the function calls and input arguments in order to use the tool. In that sense, a tool can be virtually anything, including the interface to a digital system, a scratch pad for thinking Anthropic (2025), a request to the user LangChain (2025b), another LM OpenAI (2025), or developer-defined code. Notably, the function calling corresponds to a structured output from the post-trained LM that generates function calls and input arguments. Function calling provides a principled way to use tools in the sense that it allows the output of the LM to be parsed and provides a unified way to describe available tools to the LM. Thus, we follow previous work and use the terms function calling and tool interchangeably Grattafiori et al. (2024). It is important to note that the order of arguments in tools, as well as the order of tools when calling multiple tools, matter because post-trained LMs are autoregressive models McCoy et al. (2023); Grosse et al. (2023).
In practice, a post-trained LM comes with built-in tools or developer-defined code as tools. As we will see in Sec. 3, training of tools refers to updating developer-defined code. Built-in tools OpenAI (2025); Anthropic (2025) typically includes:
@@
  • a web search API Nakano et al. (2022) allows the LM to tackle tasks with public data. The LM typically generates a response by citing the retrieved content.
@@
  • a vector database Lewis et al. (2020); Izacard & Grave (2021); Ram et al. (2023); Anthropic (2024); Jiang et al. (2020) for storing data as embeddings that can be retrieved through top-$k$ retrieval to a query. It is often used for accessing private data or as a memory module Park et al. (2023).
@@
  • a computer use tool OpenAI (2025) that interacts with a headless or headful browser to perform browsing tasks.
  • a thinking tool Anthropic (2025), unlike ReAct Yao et al. (2023b), defines reasoning itself as a callable tool, enabling structured, mid-process reflection that offers a more adaptable alternative—allowing the language model to interleave reasoning steps with multiple tool calls to maintain coherence and improve complex task handling.
@@ 2.3 Multi-Step Process
Figure 4: Multi-step process of a scaffolded LM. The distinctive trait between agent and workflow is that for each workflow step, the developer controls the system prompt and user instruction to pre-define the sub-task at that step. This makes workflows much more efficient and predictable than agents. However, workflows are engineered for specific scenarios. In this figure, we assume the case where the tool execution happens on the developer side. For built-in tools, tool execution happens on the model (API) side OpenAI (2025).
A scaffolded LM involves a multi-step process with tools, where at each step, the LM outputs tool calls. As shown in Fig. 4, guiding the model at each step with tailored prompts reduces its decision-making burden. Thus, in practice, a workflow involves pre-defining the number of steps and crafting appropriate user or system prompts for each, while an agent focuses on engineering the system prompt and the set of tools.
These approaches mainly differ in how much autonomy the LM has, and we can also refer to this autonomy as the degree of agenticness LangChain (2025a). An agent works well in unstructured scenarios, i.e., in cases where it is hard to anticipate all the scenarios that the LM might face. In contrast, a workflow approach works well when we know the solution path to a problem and can thus directly define each step required to solve it. A good example is solving GitHub issues in SWE-Bench Jimenez et al. (2024). A workflow such as Agentless Xia et al. (2024) solves GitHub issues with the following steps. Roughly, it first calls the LLM with the GitHub repository and issue description to provide a list of files that potentially contain the bug. Those files are then used to prompt the LLM to choose relevant content inside those files. Then, it asks the LLM to generate multiple patches. In comparison, an agent such as Openhands Wang et al. (2024c) provides the LM with a terminal and a code interpreter to interact with the repository.
@@ 2.3.1 Agents
Agents solve tasks by deciding at each step which tool to use and how to use it. They also determine when a task is solved based on their own interpretation of results and goals. An agent starts with a system message and a user message and generates the first response. Then, at each step, it appends to the input $\mathbf{x}$ the LM response and tool execution response, and generates the next response provided with the tool execution response. It continues until it decides to stop or it reaches the maximum number of steps, and (optionally) returns a response to the user. Note that while Fig. 3 appends the LM response and tool response under the special headers, we can also append them to the user message. Besides providing the right tool to our agent, an important design of the agent is how it generates function calls and input arguments. In particular, it affects the horizon as well as the task complexity. Traditionally, people have used structured formats such as JSON or XML to generate tool calls. Recently, CodeAct Wang et al. (2024b) proposes to use executable Python code as a unified way to call tools. As shown in Fig. 5, compared to structured formats like XML, code inherently supports both control flow (e.g., conditionals and loops) and data flow, allowing intermediate results to be stored and reused via variables when coordinating multiple tool calls. They Wang et al. (2024b) find that it achieves comparable or better performance on atomic tool use, while excelling in tasks requiring multiple calls to multiple tools in multi-turn interactions. However, it requires a sandboxed environment for code execution.
@@ 2.3.2 Workflows
Workflows Xia et al. (2024); Lu et al. (2024); Yamada et al. (2025) shine in applications where a solution path can be pre-determined. In contrast to agents, they are rendered more predictable and reliable by limiting the LM’s autonomy. At each step, the input $\mathbf{x}$ consists of a system message and a user message that is updated to specify the sub-task at each step. Thus, workflows consist of a pre-determined number of steps. It is more cost-efficient as it avoids any unnecessary exploration steps Xia et al. (2024). Unlike agents that scale with the number of steps Pan et al. (2024), workflows scale by generating multiple responses in parallel Snell et al. (2024); Welleck et al. (2024); Chen et al. (2024a). Specifically, by fixing the solution path, the output format at each step is predictable. Thus, workflows use meta-generation techniques Welleck et al. (2024) such as best of $n$ sampling or majority voting Wang et al. (2023b) to scale test-time compute. For instance, in Yamada et al. (2025), the AI scientist-v2 workflow can reach workshop-level papers by implementing tree-based methods on top of its previous version Lu et al. (2024). In particular, by generating multiple solutions at each step, it effectively generates multiple solution paths, and then uses heuristics (e.g., LLM as judge Yao et al. (2023a)) to select the best-performing path.
Figure 5: Generating function calls and input arguments. We show an agent that uses a tool with structured format or executable code Wang et al. (2024b). Provided with a system prompt to control how it interacts with the tool, the model either generates output that is easily parsed or directly generates Python code that executes the tool.
@@ 2.4 Benchmarks
As scaffolded language models (LMs) become increasingly integrated into real-world Windsurf ; Wang et al. (2024c); Cursor (2023); Microsoft (2025), there is growing momentum toward developing benchmarks that reflect mixed autonomy settings—scenarios where humans and AI share control and collaborate to achieve shared goals. In practice, these models are not isolated problem-solvers; they operate through the same APIs, UIs, and tools as human experts, and seldom succeed on the first try Rein et al. (2025); Wijk et al. (2024); Wei et al. (2025a); Yoran et al. (2024); Jimenez et al. (2024); Pan et al. (2024); Zhong et al. (2025); Jha et al. (2025); Starace et al. (2025). This necessitates multi-turn interactions, human-in-the-loop feedback, and iterative refinement—dynamics absent from traditional benchmarks focused on raw intelligence (e.g., math Olympiads or code generation).
To evaluate real-world effectiveness, recent benchmarks have shifted from measuring standalone task accuracy to assessing human-AI collaboration and productivity gains. For example, Tau-Bench Yao et al. (2024a) and Sweet-RL Zhou et al. (2025) simulate human feedback using LMs to better reflect the back-and-forth nature of co-working. Benchmarks like HCAST Rein et al. (2025) and RE-Bench Wijk et al. (2024) calibrate task difficulty against human effort—reporting metrics like “fraction of human pace” rather than raw accuracy. SWE-Bench Jimenez et al. (2024) and SWE-Lancer Miserendino et al. (2025) go further, grounding evaluation in software engineering tasks from GitHub or Upwork—tasks that are compensated in real-world settings and evaluated via functional test cases.
This benchmarking evolution reflects a shift: from measuring intelligence to measuring impact OpenAI (2018). The goal is no longer to assess whether a model solves a task perfectly in isolation, but how reliably and usefully it supports human workflows. Mixed autonomy systems must fluidly ratchet delegation up or down, adapt to under-specified tasks, navigate unclear success criteria, and coordinate across asynchronous streams of work. As highlighted by the METR report, such messiness is correlated with task length and reflects the realities of intellectual labor. Accordingly, we argue for new evaluation paradigms that go beyond single-turn metrics and instead quantify collaborative efficiency, reliability, and human-labor savings in dynamic, real-world environments.
@@ 3 Training with Language Supervision
We view scaffolded LM as semi-parametric models where the LM is the parametric component $\theta$ that is fixed. Instead, we consider the training of non-parametric variables $\phi$ that are represented in language or code. Thus, we denote our scaffolded LM as $\pi_{\phi}$, where we omit $\theta$ as it is fixed. As shown in Fig. 7, non-parametric variables $\phi$ include the system prompt, the user input template, developer-defined code as tools, or the scaffold’s code. As shown in Fig. 6, we use an LM as an optimizer to generate the updated variables $\phi$ according to an objective $J(\phi)$, an execution trace $\tau$, and optionally execution feedback $f$, all provided in natural language.
As in machine learning (ML), we assume a training, validation, and test dataset sampled from the same (unknown) distribution. During training, we iteratively sample data point(s) and update $\phi$. At inference, $\phi$ is fixed. Unlike in ML, where we train with Monte Carlo samples to prevent catastrophic forgetting Hadsell et al. (2020), an LM optimizer works with interpretable variables $\phi$ where the role of batching is still an open research question Yuksekgonul et al. (2025). The datasets $\mathcal{D}$ consist of natural language queries $\{\mathbf{q}\}$ corresponding to user input from which we obtain execution traces $\tau\sim\pi_{\phi}(\mathbf{q})$. During training, the execution traces come from any scaffolded LM $\pi_{\phi^{\prime}}$ or the current scaffolded LM $\pi_{\phi}$. In that sense, it shares similarity with RL where we have on/off-policy training, offline/online training Sutton & Barto (2018). However, the training dynamics significantly differ from those in RL. While RL is about finding good data (exploration) and imitating that good data (exploitation) Eysenbach et al. (2020), recent works in Sec. 3.3 learn only from bad execution traces. This is because failed execution traces often yield rich textual feedback, such as error messages or simulated user responses.
@@
Central to this paradigm is an optimizer implemented with an LM as opposed to bootstrapping successful execution traces as in-context learning examples Khattab et al. (2022); Von Oswald et al. (2023). We further discuss how this paradigm differs from other approaches that have employed an LM and non-parametric components for improving a scaffold LM. First, in Reflexion Shinn et al. (2023), the updates are not retrained across tasks and are only used for retrying a task. Second, CoALA Sumers et al. (2024) is an orthogonal approach that works at run-time with a specific scaffold. In particular, it requires changing the design of the agent with a memory module. This memory module not only serves as long-context but also for improving the agent over time. However, this requires the agent to make additional decisions on what and when to store or retrieve knowledge from this memory module. In training with language supervision, recent works focus on a setting where we are provided with a fixed dataset for training. Thus, it can be seen as amortizing the expensive inference of cognitive architecture Sumers et al. (2024) into a training procedure.

### v3

@@ 3.1 Post-trained Language Models
\__box_backend_scale:Nnn System You have access to the following functions. Respond in JSON format. { "name": "search_web", "description": "Search the web for current information", "parameters": { "type": "object", "properties": { "query": { "type": "string", "description": "The search query" } }, "required": ["query"] } } You are an assistant that provides the user with information from the web. User Find the latest trip of IShowSpeed Assistant Tool Response IShowSpeed’s most recent trip was a high-profile tour across China and Mongolia in March and April 2025.
Figure 4: Example of chat history in LLaMA-3. Adapted from [61]. A chat history starts with a prompt comprising a system and user message in LLaMA-3. The system message enumerates available tools (as JSON schemas) and sets high-level behavior. The user message specifies the task and the data. During the multi-step process, the chat history is updated with assistant messages (LM responses), user feedback, and tool responses.
From a probabilistic perspective, given an input (corresponding to the chat history) $\mathbf{x}=(x_{1},\dots,x_{T})$ consisting of a sequence of tokens $x_{t}\in\mathcal{V}$, a post-trained LM generates a response $\mathbf{y}=(y_{1},\dots,y_{N})$ conditioned on $\mathbf{x}$ in an autoregressive manner:
@@
During post-training, the LM is trained to follow a chat format. Thus, the response follows the chat format and corresponds to a tool call or a response to the user. As shown in Fig. 4, the multi-step process in a scaffolded LM produces a chat history where messages come from different roles. Specifically, a chat format is a protocol that specifies how to convert a chat history into a single sequence of tokens $\mathbf{x}$ that the post-trained LM both expects as input and emits as output. For instance, the “tool” role is used for tool responses. In other words, we can understand a chat format as a protocol consisting of
  • 1.
    a finite set of roles, e.g., system, user, assistant, tool.
  • 2.
    a set of special tokens that act as markers, such as role headers, channel indicators, message delimiters, tool-call fences, and termination markers [62].
Importantly, the system message and user message specify how the model should behave as well as when and how to use tools. A chat starts with a system message and a user message, and we refer to them as the prompt. Non-parametric training steers the behavior of scaffolded LMs by optimizing the prompt and tools.
By default, it is recommended to adhere to the provider’s native chat format with explicit role headers (system, user, assistant, tool), placing tool outputs under the dedicated tool role and assistant generations under the assistant role. This aligns with how post-trained LMs are supervised and preserves the intended privilege ordering [63]. During post-training, the LM is explicitly trained to treat each message source (system → user → tool output) as having an explicit privilege order [63]. Thus, if there’s a conflict between messages from different sources, the LM is trained to prioritize them based on this privilege order. This avoids prompt-injection [64], jailbreaks [65], and system-prompt extraction [66]. This process is different from knowledge conflict [67, 68] that focuses on factuality. It arises when there’s a conflict between parametric knowledge (what the model “knows”), contextual knowledge (retrieved content), and instruction knowledge (user/system assumptions or edits). Accordingly, Claude’s system prompt includes explicit instructions to avoid invoking external tools on topics the model already “knows,” which could be a strategy to avoid knowledge conflicts to some extent [66, 69].
@@ 3.2 Tools
A tool can be virtually anything, including the interface to a digital system, a scratch pad for thinking [70], a request to the user [71], another LM [72], or developer-defined code. We define a tool as an external capability that the LM can invoke via the tool API [7, 8]. Moreover, we define a function as a tool defined by a JSON schema. Similarly, we define tool calling as a specific type of response from the LM that includes a call to one of the available tools and a corresponding plain text for that tool (e.g., a query to a search engine). We define function calling as a specific type of tool call that corresponds to a structured JSON. It is important to note that the order of arguments in a function matters because post-trained LMs are autoregressive models [73, 74].
During post-training, LMs may be trained to use built-in tools, which we call tool-integrated reasoning [75]. Built-in tools are executed remotely on the provider side within temporary secure sandboxes, rather than on the user’s local machine. Thus, built-in tools are only used for seeking external information to improve their reasoning process rather than advancing on the task itself. Built-in tools [23, 70] typically include:
@@
  • a web search API [76] allows the LM to tackle tasks with public data. The LM typically generates a response by citing the retrieved content.
@@
  • a vector database [77, 78, 79, 80, 81] for storing data as embeddings that can be retrieved through top-$k$ retrieval for a query. It is often used for accessing private data or as a memory module [82].
@@
  • a computer use tool [83] that interacts with a headful browser. Unlike the web search API, it allows the LM to use interactive websites.
The multi-step process in scaffolded LMs naturally allows sequential tool composition, where tools are called sequentially in a coordinated manner to solve a task. In that sense, scaffolded LMs start by producing a plan in plain text to improve tool composition [75]. CodeAct [84] proposes to use executable Python code as a unified way to call tools. As shown in Fig. 6, code inherently supports both control flow (e.g., conditionals and loops) and data flow, allowing intermediate results to be stored and reused via variables when coordinating multiple tool calls. They [84] find that it achieves comparable or better performance on sequential tool composition. However, this approach is vulnerable to executing unsafe code and requires sandboxing. Even with sandboxes, strict capability scoping and comprehensive controls are needed: network and file-system egress restrictions, timeouts, resource quotas, import allow-lists, AST/static analysis, and immutable audit trails. Building and operating that infrastructure is non-trivial. Moreover, code-injection attacks become more dangerous because untrusted tool outputs (e.g., web pages or file contents) can be stitched into code and executed. Additionally, agents with access to the internet are exposed to a wide variety of attacks where tools return instructions, when followed by the LM, leads to leaking personal data [64].
Besides, LMs may also occasionally generate invalid tool calls (e.g., syntax errors), wrong tools, or irrelevant tools. For instance, irrelevant tool calls may be a result of reward hacking during post-training [85]. Similarly, the LM may face unexpected tool execution failures (e.g., rate-limit). To account for these, developers must implement error-handling mechanisms to catch exceptions during tool execution and provide feedback to the LM for recovery. Additionally, the way tools are defined and described inside the system message affects reliability. Clear, specific descriptions and intuitive names help the LM choose the right tool. Additionally, clear description provides explicit guidance for the LM to learn when to use this tool as opposed to learning it through RL which is prone to reward hacking [85]. Thus, the optimization of prompts and tools is important.
@@ 3.3 Multi-Step Process
Figure 5: Multi-step process of a scaffolded LM. The distinctive trait between agent and workflow is that for each workflow step, a different prompt is written to pre-define the sub-task at that step. This makes workflows much more efficient and predictable than agents. However, workflows are engineered for specific scenarios. In this figure, we assume the case where the tool execution happens on the developer side. For built-in tools, tool execution happens on the provider’s side [23].
A scaffolded LM involves a multi-step process with tools. Each step consists of an interaction between the assistant and tool. As shown in Fig. 5, guiding the model at each step with tailored prompts reduces its decision-making burden. Thus, in practice, a workflow involves pre-defining the number of steps and crafting appropriate prompt for each step, while an agent focuses on engineering a single prompt and the set of tools.
These approaches mainly differ in how much autonomy the LM has, and we can also refer to this autonomy as the degree of agenticness [86]. An agent works well in unstructured scenarios, i.e., in cases where it is hard to anticipate all the scenarios that the LM might face. In contrast, a workflow approach works well when we know the solution path to a problem and can thus directly define each step required to solve it. A good example is solving GitHub issues in SWE-Bench [50]. A workflow such as Agentless [49] solves GitHub issues with the following steps. Roughly, it first calls the LLM with the GitHub repository and issue description to provide a list of files that potentially contain the bug. Those files are then used to prompt the LLM to choose relevant content inside those files. Then, it asks the LLM to generate multiple patches. In comparison, an agent such as Openhands [5] provides the LM with a terminal and a code interpreter to interact with the repository.
@@ 3.3.1 Agents
Agents solve tasks by deciding at each step which tool to use and how to use it. They also determine when a task is solved based on their own interpretation of results and goals. The chat of an agent starts with a prompt and each step updates the chat history. The agent continues until it decides to stop or it reaches the maximum number of steps, and returns a response to the user that summarize what it did. Besides providing the right tool to our agent, an important design of the agent is how it composes tools. In particular, code execution [84] offers a unified way to compose tools with control flow and data flow, and typically results in a shorter horizon. Another approach prompts the agent to provide a plan beforehand [75].
@@ 3.3.2 Workflows
Workflows [49, 87, 88] shine in applications where a solution path can be pre-determined. In contrast to agents, they are rendered more predictable and reliable by limiting the LM’s autonomy. At each step, the input $\mathbf{x}$ consists of a different prompt that specify the sub-task at that step. Thus, workflows consist of a pre-determined number of steps. It is more cost-efficient as it avoids any unnecessary exploration steps [49]. Unlike agents that scale with the number of steps [89], workflows scale by generating multiple responses in parallel [90, 56, 91]. Specifically, by fixing the solution path, the output format at each step is predictable. Thus, workflows use meta-generation techniques [56] such as best of $n$ sampling or majority voting [92] to scale test-time compute. For instance, in [88], the AI scientist-v2 workflow can reach workshop-level papers by implementing tree-based methods on top of its previous version [87]. In particular, by generating multiple solutions at each step, it effectively generates multiple solution paths, and then uses heuristics (e.g., LLM as judge [93]) to select the best-performing path. The cost of parallelization in workflows mainly depends on the aggregation method. For instance LLM as judge adds one extra LM call, while pairwise judging adds quadratic cost. Meanwhile, aggregation can be done without an LM judge, e.g., unit tests [49] or majority voting [92].
Figure 6: Tool compositionality. Left: tools are composed sequentially. As shown in the figure, some LMs support generating independent tool calls in parallel [7]. Right: executable Python code as a unified way to call tools [84]. Code inherently supports both control flow (e.g., conditionals and loops) and data flow, allowing intermediate results to be stored and reused via variables when coordinating multiple tool calls
@@ 3.4 Benchmarks
We highlight a growing trend in recent benchmarks: from measuring pure intelligence (e.g., math olympiads) to measuring productivity gains (e.g., software engineering) [4]. The goal is no longer to assess whether a model solves a task perfectly in isolation, but how reliably and usefully it supports human workflows. Importantly, these benchmarks are used to compare LMs on products [94, 5, 95, 96] that operate in mixed-autonomy. This is a setting where humans and AI work together, sharing control and decision-making responsibilities to achieve a goal. For instance, in vibe coding, the human guides intent and judges quality, while the AI proposes implementations and refactors.
To accurately gauge mixed-autonomy capability, these benchmarks align agent performance with human effort—measuring monetary return [97], simulating human collaborators [98, 99], exercising real software and APIs [50, 100, 101], and using time as a measure of task difficulty [102] because time is money.
SWE-bench [50] and SWE-Lancer [97] ground evaluation in software engineering tasks from GitHub or Upwork—tasks that are compensated in real-world settings and evaluated via functional test cases. Other benchmarks go beyond measuring raw accuracy. BrowseComp [100] evaluates accuracy and calibration error. In particular, the LM is asked to output a confidence (0–100) with its answer; BrowseComp then evaluates how well that confidence aligns with reality. HCAST [102] evaluates success rate as a function of human-calibrated task length (e.g., <15min, 15min-15h, 1-4h, 4h+). Additionally, they compute a “50% time-horizon”: the human task duration at which the agent has a 50% chance of success, fitted from per-task success vs. human time. RE-Bench [103] finds that currently humans scale better with increased time budgets to solve the task. For example, $\tau$-Bench [98] and Sweet-RL [99] simulate human feedback using LMs to better reflect the back-and-forth nature of co-working. For tasks that are hard to verify, PaperBench [101] proposes a rubric organized as a hierarchical tree of requirements. The leaf nodes are precise, binary checks (pass/fail). Internal nodes aggregate their children by a manual, importance-based weighting, and the root score is the paper’s Replication Score. Tab. 1 summarizes these benchmarks.
Benchmark Scenario Evaluation metrics SWE-Bench [50] Software engineering Unit test. SWE-Lancer [97] Software engineering pass@1 / pass@k, dollars earned, unit test, SWE Manager by matching the original manager’s proposal. HCAST [102] Machine learning, software engineering, cybersecurity Binarized success rate by human-time buckets. RE-Bench [103] AI Research Expert-graded task success; fraction of human pace; time-to-solution. PaperBench [101] AI research Rubric organized as a hierarchical tree of requirements. $\tau$-Bench [98] Tool-agent-user interaction pass@k metric as a measure of reliability of an agent. SWEET-RL [99] Multi-turn collaborative reasoning Task success/win rate under simulated feedback. BrowseComp [100] Deep research Accuracy, calibration error.
Table 1: Benchmarks commonly used for scaffolded LMs.
@@ 4 Training with Language Supervision
We view scaffolded LM $\pi$ as semi-parametric models where the LM is the parametric component that is fixed. Instead, we consider the training of non-parametric variables, prompt and tools, that are represented in language and code. As shown in Fig. 8, non-parametric variables include the prompt and developer-defined code as tools. As shown in Fig. 7, we use an LM as an optimizer to generate the updated prompt or tool according to an objective (“improve prompt according to feedback on this execution trace”), an execution trace, and optionally execution feedback $f$, all provided in natural language.
As in machine learning (ML), we assume a training, validation, and test dataset sampled from the same (unknown) distribution. During training, we iteratively sample data point(s) and update prompt or tool. At inference, prompt(s) and tools are fixed (in the case of workflows there’s a multiple prompts, i.e., a different prompt for each step). Unlike in ML, where we train with Monte Carlo samples to prevent catastrophic forgetting [3], an LM optimizer works with variables represented in language or code as opposed to tensors. Thus, the role of batching is still an open research question [9]. The datasets $\mathcal{D}$ consist of user message corresponding to different tasks from which we obtain execution traces $\tau\sim\pi$. During training, the execution traces come from any scaffolded LM $\pi^{\prime}$ or the current scaffolded LM $\pi$. In that sense, it shares similarities with RL, where we have on/off-policy training and offline/online training [104]. However, the training dynamics significantly differ from those in RL. While RL is about finding good data (exploration) and imitating that good data (exploitation) [105], recent works in Sec. 4.3 learn only from bad execution traces. This is because failed execution traces often yield rich textual feedback, such as error messages or simulated user responses.
@@
Central to this paradigm is an optimizer implemented with an LM as opposed to bootstrapping successful execution traces as in-context learning examples [106, 107]. We further discuss how this paradigm differs from other approaches that have employed an LM and non-parametric components for improving a scaffolded LM. First, in Reflexion [108], the updates are not retrained across tasks and are only used for retrying a task. Second, CoALA [109] is an orthogonal approach that works at run-time with a specific scaffold. In particular, it requires changing the design of the agent with a memory module. This memory module not only serves as long-context memory but also for improving the agent over time. However, this requires the agent to make additional decisions on what and when to store or retrieve knowledge from this memory module. In training with language supervision, recent works focus on a setting where we are provided with a fixed dataset for training. Thus, it can be seen as amortizing the expensive inference of a cognitive architecture [109] into a training procedure.